X-25 Item is an important part used in the production of Laptops.

Daily use of Item X-25 in a high-tech Manufacturing Process during the Delivery or Lead time for Item X-25 is 4 working days is given below:

1) Avg Daily Use = 200 2) Maximum Daily Use = 230 3) Minimum Daily Use = 160 and SD = +/- 20

SD = Standard Deviation

Calculate the Reorder Point (ROP) for X-25 with the Safety Stock (SS) at 97% Service Level in order to prevent Stock-out. In other word Stock-out at 3% level is acceptable.

a.

920

b.
c.

426

d.

559

e.

838

f.

800
Answer #1

Reorder point = dL + zσd√L

where, where d = average daily demand = 200

σd = standard deviation of daily demand = 20

L = Lead Time = 4 days

z = number of standard deviations based on service level 97% = 1.88 from z table

Reorder Point = 200*4 + 1.88*20*√4 = 875.2